Walt made an extra $12,000 last year from a part time job. He invested part of the money at 7% and the rest at 9%. He made a total of $970 in interest. How much was invested at 7%?
Answer by MissionPossible(18)

He invested part of the money at 7% and the rest at 9%.
How much was invested at 7%? -----> X
invested part of the money at 7% would made X*0.07
invested part of the money at 9% would made (12,000 - X)*0.09
He made a total of $970 ---> 970 = X*0.07 + (12,000 - X)*0.07
970 = X*0.07 + (12,000 - X)*0.09
970 = X*0.07 + 1080 - 0.09X
0.09X-0.07X = 1080 - 970
0.02X = 110
X = 110/0.02
= 5500
How much was invested at 7%? $5,500 Ans.